OpenStudy (anonymous):

A certain strain of bacteria that is growing on your kitchen counter doubles every 15 minutes. Assuming that you start with only one bacterium, how many bacteria could be present at the end of 7 hours?

OpenStudy (amistre64):

A=Pert

OpenStudy (anonymous):

2^28

OpenStudy (amistre64):

r = ln(2)/t r = ln(2)/(1/4) = 4 ln(2) = ln(16) A = 1*e^(ln(16)*7) = 268,435,456 maybe

OpenStudy (anonymous):

268435456

OpenStudy (anonymous):

yeah.. right

OpenStudy (amistre64):

give or take a few due to evolution lol